# Exploring Imbalanced Annotations for Effective In-Context Learning

### v1

@@ Abstract
Large language models (LLMs) have shown impressive performance on downstream tasks through in-context learning (ICL), which heavily relies on the demonstrations selected from annotated datasets. Existing selection methods may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios. In this work, we show that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks and selection methods. Moreover, traditional rebalance methods fail to ameliorate the issue of class imbalance in ICL. Our method is motivated by decomposing the distributional differences between annotated and test datasets into two-component weights: class-wise weights and conditional bias. The key idea behind our method is to estimate the conditional bias by minimizing the empirical error on a balanced validation dataset and to employ the two-component weights to modify the original scoring functions during selection. Our approach can prevent selecting too many demonstrations from a single class while preserving the effectiveness of the original selection methods. Extensive experiments demonstrate the effectiveness of our method, improving the average accuracy by up to 5.46 on common benchmarks with imbalanced datasets.
Machine Learning, ICML
@@ 1 Introduction
In-context learning (ICL) has been a remarkable capability of Large Language Models (LLMs), enabling them to perform downstream tasks by simply conditioning on a few task demonstrations. (Brown et al., 2020). ICL consistently outperforms zero-shot inference across various tasks without needing parameter updates, positioning it as a strong alternative to supervised fine-tuning (SFT) (Mosbach et al., 2023; Panwar et al., 2024). In particular, previous studies typically focus on designing methods for demonstration selection from high-quality annotated datasets (Liu et al., 2022; Baldassini et al., 2024; Wang et al., 2024). The effectiveness of these selection algorithms may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios (Wei et al., 2022; De Alvis & Seneviratne, 2024) — certain classes are underrepresented (see Figure 1). Thus, it is of great importance to investigate how such imbalances affect ICL’s performance.
Figure 1: An example of imbalanced dataset. In Emotion (Saravia et al., 2018), a few sentiments make a large contribution and data tend to show a long-tailed distribution. For example, “joy“ and “sadness“ are head classes, while most other classes, such as “love“ and “surprise“, are tail classes.
In this work, we present the first study on the impact of long-tailed distribution in annotated datasets on in-context learning. First, we theoretically show that the class prior distribution affects the prediction of ICL. Empirically, we also find that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks and selection methods (refer to Figure LABEL:figure:2). Moreover, traditional rebalance methods not only fail to ameliorate but, in some cases, exacerbate the issue of class imbalance in ICL. This motivates us to develop a universal method that consistently improves ICL’s performance in the presence of imbalanced annotated datasets.
In this paper, we show that the issue of imbalanced annotated datasets can be mitigated by applying class weighting to the original scoring functions during selection. Our method is motivated by accounting for the differences between annotated and test distributions using two-component weights: class-wise weights, which assesses the difference between class prior distributions, and conditional bias, which measures the bias of class-wise weights due to different conditional distributions. Therefore, our key idea is to employ the two-component weights to rectify the original scoring functions during selection. Specifically, we employ effective number (Cui et al., 2019) as class-wise weights and estimate conditional bias by minimizing the empirical error of a balanced validation set. We then employ the two-component weights to modify the original scoring functions during selection. In effect, our approach can prevent selecting too many demonstrations from a single class while preserving the effectiveness of the original selection methods.
@@
    We find that imbalanced class distributions in annotated datasets significantly degrade the ICL’s performance. We also show that classical rebalance methods fail to ameliorate the imbalanced problems of ICL.
@@
    We propose a simple and effective method to enhance ICL’s performance with imbalanced annotated datasets. The key idea is to estimate the conditional bias and employ the two-component weights to modify the original scoring functions during selection.
@@
    We empirically show that our method can improve the performance of existing selection methods across various imbalance ratios. Meanwhile, we also validate the effectiveness of our method on generation tasks.
@@ 2.1 In-context learning
In the context of large language models (LLMs), in-context learning (ICL) aims to generate text outputs $\mathrm{\mathbf{y}}=(y_{1},...,y_{|\mathrm{\mathbf{y}}|})$ (i.e., token sequences) conditioned on input $\mathrm{\mathbf{x}}=(x_{1},...,x_{|\mathrm{\mathbf{x}}|})$ and context $\mathrm{\mathbf{C}}_{K}$. In particular, the context $\mathrm{\mathbf{C}}_{K}=\{(\mathrm{\mathbf{x}}_{i},\mathrm{\mathbf{y}}_{i})\}^{K}_{i=1}$ comprises $K$ task demonstrations (e.g. input-output pairs) selected from a large annotated dataset with $N$ examples $\mathcal{D}_{c}=\{(\mathrm{\mathbf{x}}_{i},\mathrm{\mathbf{y}}_{i})\}^{N}_{i=1}$. Let $f_{\theta}(\mathrm{\mathbf{C}}_{K},\cdot)$ be the ICL model with demonstrations $\mathrm{\mathbf{C}}_{K}$, using the LLM $f$ parameterized by $\boldsymbol{\theta}$. Given a test input $\mathrm{\mathbf{x}}_{t}$, we generate the output $\mathrm{\mathbf{y}}_{t}$ via the ICL model as:
@@
To improve the performance of ICL, previous studies (Liu et al., 2022; Peng et al., 2024; Rubin et al., 2022; Ye et al., 2023a) designed various scoring functions $s(\cdot,\cdot)$ to select demonstrations $\mathrm{\mathbf{C}}_{K}$ from an annotated dataset $\mathcal{D}_{c}$ as:
@@
While current selection methods showcase promising performance on common-used benchmarks, their effectiveness may hinge on the distribution of annotation datasets $\mathcal{D}_{c}$. For example, ICL might struggle to make accurate predictions for underrepresented groups within these annotated datasets. Real-world datasets (see Figure 1), however, often exhibit an imbalanced distribution, with a few ‘head’ classes containing many examples and numerous ‘tail’ classes having significantly fewer examples. The concern may lead to challenges in effectively employing in-context learning in real-world applications. We proceed with a formulation of the imbalanced setting of ICL.
@@ 2.2 Setting of imbalanced ICL
In this work, we begin with the class-imbalance setting of in-context learning in classification tasks [^1] In addition to labels, the imbalance can also occur among various groups, particularly in generation tasks. We extend the imbalanced setting to generation tasks in Section 6. , where the label space $\mathcal{Y}:=\{1,\ldots,K\}$. Let $n_{j}$ denote the number of instances in class $j$, where $j\in\mathcal{Y}$. In the class-imbalanced setting, the annotated dataset $\mathcal{D}_{c}$ has an unequal distribution of instances across different classes in $\mathcal{Y}$. In particular, the class distribution is such that: $n_{j}\ll n_{K}$, for some $j,k\in\mathcal{Y}$, where $j\neq K$. We quantify the imbalance ratio as $\phi=\frac{\min_{j\in\mathcal{Y}}n_{j}}{\max_{j\in\mathcal{Y}}n_{j}}$, and a higher imbalance ratio indicates a more severe class imbalance in the dataset.
In the real world, class-imbalanced distributions are frequently observed in various datasets. For instance, in the Emotion dataset (Saravia et al., 2018), the ‘Joy’ class constitutes 33% of the data, whereas the ‘Surprise’ class makes up only 4%. The CodeSearchNet dataset (Husain et al., 2019) includes 29% JavaScript code while Ruby accounts for just 3%, demonstrating the significant imbalance issue among programming languages. Therefore, it is crucial to ensure the ICL performance across all classes under the class-imbalanced setting in $\mathcal{D}_{c}$. In what follows, we analyze the effect of class imbalance on ICL from theoretical and empirical perspectives.
@@ 3 Pilot study
3.1 The effect of imbalanced dataset
3.1.1 Theoretical Analysis
We start by presenting Remark 3.1 (Xie et al., 2022), which demonstrates that ICL allows LLMs $f$ parameterized by $\theta$ to learn the context-generated distribution from demonstrations $\mathrm{\mathbf{C}}_{K}$, such that the model’s prediction approximates the output $\mathrm{\mathbf{y}}$ generated by the data generated model $\theta^{*}$ given the test input $\mathrm{\mathbf{x}}$.
Remark 3.1.
@@ 3.1.1 Theoretical Analysis
where $P_{\theta^{*}}(\mathrm{\mathbf{x}}_{c})$ and $P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t})$ represent $P_{\theta^{*}}(X=\mathrm{\mathbf{x}}_{t})$ and $P_{\theta^{*}}(Y=\mathrm{\mathbf{y}}_{t})$, respectively.
The Eq. 3 indicates that the class prior distribution $P(Y)$ affects the prediction of ICL. in real-world tasks, the class prior of the test distribution is usually balanced (i.e., a uniform distribution over labels), while the annotated dataset exhibits a long-tailed class distribution. Formally, we have $P_{c}(Y=j)\neq P_{t}(Y=j)$ for any class $j\in\mathcal{Y}$. The predicted posterior probability in Eq. 3 becomes unreliable when there is a large discrepancy of class priors between annotated dataset $P_{c}(Y)$ and test dataset $P_{t}(Y)$.
3.1.2 Empirical Analysis
We also conduct experiments on various downstream tasks, including Amazon (Keung et al., 2020), AgNews, Yahoo and Yelp (Zhang et al., 2015), NQ (Kwiatkowski et al., 2019) and CodeSearchNet (Husain et al., 2019)). To simulate the class-imbalanced setting, we generate imbalanced datasets with a pre-defined probability (e.g., $\phi$ = 0, 10, 50, 100). We evaluate the performance of ICL on a balanced test dataset with three LLMs, including Llama3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023) and OPT-6.7B (Zhang et al., 2022a). Furthermore, we also compare the performance of ICL with various set sizes and selection methods, including Random (Min et al., 2022), TopK (Liu et al., 2022), DPP (Ye et al., 2023a), VoteK (Su et al., 2023), ConE (Peng et al., 2024), ByCS (Wang et al., 2024).
@@ 3.1.2 Empirical Analysis
Through the above analysis, we find that imbalanced datasets significantly hurt the ICL’s performance. Next, we will verify whether employing some classical rebalance methods can mitigate the issue of imbalanced datasets.
3.2 The failure of classical rebalance methods
Rebalance methods are widely used to deal with the issue of imbalanced annotated datasets. One such approach is to address the bias towards the minor class by adjusting the annotated dataset. In this section, we use four different rebalance methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005) and re-weighting (Cui et al., 2019). Specifically, given an annotated dataset with $N$ examples, we select demonstrations as follows:
@@ 3.2 The failure of classical rebalance methods
$\mathrm{\mathbf{C}}_{K}$ $=\cup_{i=1}^{k}\operatorname{top}_{\frac{K}{k}}\left(\left\{s(\mathrm{\mathbf{c}}_{j},\mathrm{\mathbf{x}}_{t})\right\}_{j=1}^{n_{i}}\right),$ (5)
@@
For the over-sampling method in Eq. 4, we select demonstrations $\mathbf{c}^{{}^{\prime}}_{i}$ from an over-sampling dataset with $N^{{}^{\prime}}$ examples where we repeat the examples of tail classes until their number matches that of head classes. The under-sampling method in Eq. 4 randomly trims examples belonging to head classes until the number of examples in head classes is equal to that of tail classes. Suppose $n_{j}$ represents the size of the class to which $i$-th example belongs, stratified sampling in Eq. 5 selects $\frac{K}{k}$ demonstrations from each class with $n_{j}$ examples. For the re-weighting method in Eq. 6, according to the classical reweighing method for fine-tuning (Cui et al., 2019), we select the top $K$ examples based on the scoring functions $s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ multiplied by the class-wise weights $w_{j}$ where $w_{j}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$.
We conduct experiments to verify the effectiveness of these rebalance strategies across various downstream tasks within the framework of ICL. The results in Figure 3 show that employing over-sampling and under-sampling can improve ICL’s performance when annotated datasets exhibit long-tail distributions. However, these rebalance strategies can only achieve limited effects for most downstream tasks, indicating poor generalization to ICL. Specifically, the limited effect of the oversampling method (Chawla et al., 2002) suggests that repeated examples do not provide additional information to LLMs to enhance the performance of ICL. Under-sampling (Liu et al., 2008) may remove key information of head classes. Notably, the failure of the stratified sampling method in ICL is mainly due to the decreased performance of selection methods such as TopK (Liu et al., 2022) and DPP (Ye et al., 2023a). The detailed results of generation tasks have been reported in Section 6.1.
Our empirical analysis shows that existing rebalance methods cannot effectively address the issue of imbalanced datasets for ICL. This motivates us to design new rebalance methods for ICL that can universally enhance the performance of existing selection methods.
Figure 3: Comparison of performances on four classification tasks with different imbalance ratios using four different rebalance methods. We report the average accuracy (%) of six existing selection methods ($K=8$) using OPT-6.7B (Zhang et al., 2022a).

### v2

@@ Abstract
Large language models (LLMs) have shown impressive performance on downstream tasks through in-context learning (ICL), which heavily relies on the demonstrations selected from annotated datasets. However, these datasets often exhibit long-tailed class distributions in real-world scenarios, leading to biased demonstration selection. In this work, we show that such class imbalances significantly degrade the ICL performance across various tasks, regardless of selection methods. Moreover, classical rebalancing methods, which focus solely on class weights, yield poor performance due to neglecting condition bias–skewed feature distributions within classes. To address this, we propose Reweighting with Conditional Bias (dubbed RCB), a simple and complementary approach to enhance ICL performance under class imbalance. In particular, RCB estimates conditional bias using a balanced subset and re-weights demonstration scores based on both class weight and conditional bias. In effect, RCB prevents over-selection from dominant classes while preserving the efficacy of current selection methods. Extensive experiments on common benchmarks demonstrate the effectiveness of our method, improving the average accuracy of current selection methods by up to 5.42%.
@@ 1 Introduction
In-context learning (ICL) has been a remarkable capability of Large Language Models (LLMs), enabling them to perform downstream tasks by simply conditioning on a few task demonstrations. [1]. ICL consistently outperforms zero-shot inference across various tasks without needing parameter updates, positioning it as a strong alternative to supervised fine-tuning (SFT) [2, 3]. Previous studies typically focus on designing methods for demonstration selection from high-quality annotated datasets [4, 5, 6]. The effectiveness of these selection algorithms may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios [7, 8] — certain classes are underrepresented (see Figure 5). Thus, it is of great importance to explore the effect of imbalanced annotations on in-context learning.
In this work, we present the first study on the impact of long-tailed distribution in annotated datasets on in-context learning. First, we theoretically show that the class prior distribution affects the prediction of ICL. Empirically, we also find that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks, regardless of selection methods (refer to Figure 1). Moreover, traditional rebalancing methods not only fail to ameliorate but, in some cases, exacerbate the issue of class imbalance in ICL. This motivates us to develop a universal method that consistently improves ICL’s performance in the presence of imbalanced annotations.
In this paper, we show that the issue of imbalanced annotations can be mitigated by reweighting the original scores (i.e., cosine similarity score [4]) of imbalanced examples during selection. Our method, Reweighting with Conditional Bias (dubbed RCB), is motivated by our analysis of decomposing the distributional differences between imbalanced and test datasets. We find that imbalanced datasets affect the performance of ICL through two lenses: class weights – the difference between class prior distributions – and conditional bias, which measures the bias of feature distributions within classes. The latter is typically neglected in existing rebalancing methods, leading to suboptimal improvements in ICL performance with imbalanced datasets.
Therefore, our key idea behind RCB is to incorporate conditional bias into demonstration reweighting during selection. Specifically, we calculate the effective number [9] as class weights and estimate conditional bias using Bayesian optimization on a balanced subset. Subsequently, a two-component weight containing both class weights and conditional bias is then used to reweight the demonstration scores for selection. In effect, our approach prevents over-selection from dominant classes while preserving the efficacy of current selection methods.
@@
    We present a phenomenon: imbalanced annotations significantly degrade the performance of ICL regardless of demonstration selection methods. Moreover, classical rebalancing methods fail to address this issue.
@@
    We propose a simple and complementary method by involving conditional bias to enhance ICL’s performance under class imbalance. Our method is computationally efficient and agnostic to demonstration selection methods.
@@
    We empirically validate that our methods can improve the ICL performance in both classification and generation tasks across various imbalance ratios. Our method can be applied to both open-weight LLMs and APIs, as it only requires access to model outputs.
@@ 2.1 In-context learning
In the context of large language models (LLMs), in-context learning (ICL) aims to generate text outputs $\mathrm{\mathbf{y}}=(y_{1},...,y_{|\mathrm{\mathbf{y}}|})$ (i.e., token sequences) conditioned on input $\mathrm{\mathbf{x}}=(x_{1},...,x_{|\mathrm{\mathbf{x}}|})$ and context $\mathrm{\mathbf{C}}_{K}$. In particular, the context $\mathrm{\mathbf{C}}_{K}=\{(\mathrm{\mathbf{x}}_{i},\mathrm{\mathbf{y}}_{i})\}^{K}_{i=1}$ comprises $K$ task demonstrations (e.g. input-output pairs) selected from a large annotated dataset with $N$ examples $\mathcal{D}_{c}=\{(\mathrm{\mathbf{x}}_{i},\mathrm{\mathbf{y}}_{i})\}^{N}_{i=1}$. Let $f_{\theta}(\mathrm{\mathbf{C}}_{K},\cdot)$ be the ICL model with demonstrations $\mathrm{\mathbf{C}}_{K}$, using the LLM $f$ parameterized by $\bm{\theta}$. Given a test input $\mathrm{\mathbf{x}}_{t}$, we generate the output $\mathrm{\mathbf{y}}_{t}$ via the ICL model as:
@@
To improve the performance of ICL, previous studies [4, 15, 16, 17] designed various scoring functions $s(\cdot,\cdot)$ to select demonstrations $\mathrm{\mathbf{C}}_{K}$ from an annotated dataset $\mathcal{D}_{c}$ as:
@@
While current selection methods showcase promising performance on commonly used benchmarks, their effectiveness may hinge on the distribution of annotation datasets $\mathcal{D}_{c}$. For example, ICL might struggle to make accurate predictions for underrepresented groups within these annotated datasets. Real-world datasets (see Figure 5), however, often exhibit an imbalanced distribution, with a few ‘head’ classes containing many examples and numerous ‘tail’ classes having significantly fewer examples. The concern may lead to challenges in effectively employing in-context learning in real-world applications. We proceed with a formulation of the imbalanced setting of ICL.
@@ 2.2 Setting of imbalanced ICL
In this work, we begin with the class-imbalance setting of in-context learning in classification tasks [^1] In addition to labels, the imbalance can also occur among various groups, particularly in generation tasks. We extend the imbalanced setting to generation tasks in Section 6.1. , where the label space $\mathcal{Y}:=\{1,\ldots,k\}$. Let $n_{j}$ denote the number of instances in class $j$, where $j\in\mathcal{Y}$. In the class-imbalanced setting, the annotated dataset $\mathcal{D}_{c}$ has an unequal distribution of instances across different classes in $\mathcal{Y}$. In particular, the class distribution is such that: $n_{j}\ll n_{k}$, for some $j,k\in\mathcal{Y}$, where $j\neq k$. We quantify the imbalance ratio as $\phi=\frac{\max_{j\in\mathcal{Y}}n_{j}}{\min_{j\in\mathcal{Y}}n_{j}}$, and a higher imbalance ratio indicates a more severe class imbalance in the dataset.
In the real world, class-imbalanced distributions are frequently observed in various datasets. For instance, in the Emotion dataset [12], the ‘Joy’ class constitutes 33% of the data, whereas the ‘Surprise’ class makes up only 4%. The CodeSearchNet dataset [14] includes 29% JavaScript while Ruby accounts for just 3%, demonstrating the significant imbalance issue. Therefore, it is crucial to ensure the ICL performance across all classes under the class-imbalanced setting in $\mathcal{D}_{c}$. In what follows, we analyze the effect of class imbalance on ICL from theoretical and empirical perspectives.
@@ 3 Pilot study
3.1 The effect of imbalanced data
Theoretical Analysis
In this section, we start by presenting Remark 1 [18], which demonstrates that ICL allows LLMs $f$ parameterized by $\theta$ to learn the context-generated distribution from demonstrations $\mathrm{\mathbf{C}}_{K}$, such that the model’s prediction approximates the output $\mathrm{\mathbf{y}}$ generated by the data-generating model $\theta^{*}$ given the test input $\mathrm{\mathbf{x}}$.
Remark 1
@@ 3.1 The effect of imbalanced data
where $P_{\theta^{*}}(\mathrm{\mathbf{x}}_{t})$ and $P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t})$ represent $P_{\theta^{*}}(X=\mathrm{\mathbf{x}}_{t})$ and $P_{\theta^{*}}(Y=\mathrm{\mathbf{y}}_{t})$, respectively. The Eq.  (3) indicates that the class prior distribution $P(Y)$ affects the prediction of ICL. For the class-imbalanced setting, the class prior of the test distribution is usually balanced, whereas the annotated dataset exhibits a long-tailed class distribution. The predicted posterior probability in Eq.  (3) becomes unreliable when we have $P_{c}(Y=j)\neq P_{t}(Y=j)$ for any class $j\in\mathcal{Y}$. Overall, the above theoretical analysis shows that class-imbalanced annotations cause a mismatch between the class prior distributions of the annotated and test datasets, leading to unreliable posterior probability predictions in ICL. Our experimental results further validate the analysis.
Empirical Analysis
Figure 1: The impact of imbalanced annotations on ICL performance. (a) Overall accuracy across four classification tasks. (b) Average accuracy for each class in AgNews. (c) Overall accuracy across four classification tasks using classical rebalancing methods.
We also conduct experiments on various downstream tasks, including Amazon [10], AgNews, Yahoo, Yelp [11], NQ [13], and CodeSearchNet [14]). To simulate the class-imbalanced setting, we generate imbalanced datasets with a pre-defined probability (e.g., $\phi$ = 1, 10, 50, 100). We evaluate the performance of ICL on a balanced test dataset with various LLMs, including OPT-6.7B, -13B, -30B [19], LLAMA-3-8B and -70B [20]; and APIs: ChatGPT-3.5-Turbo [21] and Gemini-2.0-Flash [22]. We also investigate the performance of ICL with imbalanced annotations using various sizes of demonstrations and selection methods.
@@
3.2 The failure of classical rebalancing methods
One simple and intuitive approach to deal with the class-imbalanced problem is rebalancing [24]. In this section, we examine whether four classical rebalancing methods—over-sampling [25], under-sampling [26], stratified sampling [27], and re-weighting [9]—can mitigate the negative effects of imbalanced annotations on ICL. Specifically, we select top K examples ranked by the score function $s(\cdot,\cdot)$ from an annotated dataset with $N$ examples as demonstrations:
@@ 3.2 The failure of classical rebalancing methods
$\mathrm{\mathbf{C}}_{K}$ $=\sum_{i=1}^{k}\operatorname{top}_{\frac{K}{k}}\left(\left\{s(\mathrm{\mathbf{c}}_{j},\mathrm{\mathbf{x}}_{t})\right\}_{j=1}^{n_{i}}\right),$ (5)
@@
For the over-sampling method in Eq. (4), we select demonstrations $\mathbf{c}^{{}^{\prime}}_{i}$ from an over-sampling dataset with $N^{{}^{\prime}}$ examples, where we repeat the examples of tail classes until their number matches that of head classes. The under-sampling method in Eq. (4) randomly trims examples belonging to head classes until the number of examples in head classes is equal to that of tail classes. Suppose $n_{j}$ represents the size of the class to which the $i$-th example belongs, stratified sampling in Eq. (5) selects $\frac{K}{k}$ demonstrations from each class with $n_{j}$ examples. For the re-weighting method in Eq. (6), we select the top $K$ examples based on the scoring functions $s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ multiplied by the class weights $w_{j}$ where $w_{j}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$ [9].
Figure 1 (c) shows that the rebalancing methods can only achieve limited effects for most downstream tasks, indicating poor generalization to ICL. Intuitively, oversampling uses repeated examples that may not provide additional information to LLMs for ICL performance, while undersampling may remove key information of head classes. For instance, with an imbalanced ratio of 100, over-sampling boosts the performance of ICL from 45.10 to 46.30, yielding limited improvement. In contrast, stratified sampling, which selects demonstrations equally from each class, reduces the performance of ICL from 45.10 to 42.46, as it undermines the effectiveness of high-performing selection methods like TopK [4]. Notably, we provide an in-depth understanding of the failure through the decomposition of the distributional differences in Subsection 4.1.